How many pounds of an $83 per pound substance must be mixed with an $11 per pound substance to get 240 lbs of an alloy which will cost $28 per pound? Round to the nearest pound.
Answer by stanbon(75887)

How many pounds of an $83 per pound substance must be mixed with an $11 per pound substance to get 240 lbs of an alloy which will cost $28 per pound? Round to the nearest pound.
---
Equation:
active + active = active
83x + 11(240-x) = 28*240
---
83x + 11*240 - 11x = 28*240
72x = 17*240
x = 56 2/3 lb. (amt. of $83 substance needed)
---
240-x = 183 1/3 lb (amt. of $11 substance needed)
================
Cheers,
Stan H.
================